Desjarlais Corporation uses the following activity rates from its activity-based costing to assign overhead costs to products. (Do not round intermediate calculations. Round your final answers to 2 decimal places.)

Activity Cost Pools
Activity Rate
Setting up batches $

169.55

per batch
Assembling products $

3.58

per assembly hour
Processing customer orders $

52.83

per customer order

Data concerning two products appear below:

Product S96U Product Q06F
Number of batches

40

46
Number of assembly hours 133 908
Number of customer orders 17 43

Required:

a. How much overhead cost would be assigned to Product S96U using the company's activity-based costing system?

b. How much overhead cost would be assigned to Product Q06F using the company's activity-based costing system?
Answer #1

Answer

--Requirements, with calculations

Overhead cost assigned: Product S96U Product Q06F
Setting up batches $6,782.00 $7,799.30
Assembling products $476.14 $3,250.64
Processing customer orders $898.11 $2,271.69
Total Overhead cost assingned $8,156.25 $13,321.63
Answer to Requirement [a] Answer to Requirement [a]

--Calculation

Overhead cost assigned: Product S96U Product Q06F
Setting up batches =40*169.55 =46*169.55
Assembling products =133*3.58 =908*3.58
Processing customer orders =17*52.83 =43*52.83